CH 24 Q 30 GoA stationary particle of charge q = 2.3 × 10-8 C is placed in a laser beam (an electromagnetic wave) whose intensity is 2.6 × 103 W/m2. Determine the maximum magnitude of the (a) electric and (b) magnetic forces exerted on the charge. If the charge is moving at a speed of 3.7 × 104 m/s perpendicular to the magnetic field of the electromagnetic wave, find the maximum magnitudes of the (c) electric and (d) magnetic forces exerted on the particle.
Answer #1

a)

I = 0.5 E^2 / ( c uo)

2.6* 10^3 = 0.5* E^2 / (3* 10^8* 4* 3.14* 10^-7)

E = 1399.77 N/C

force experienced by charge

F = qE = 2.3* 10^-8* 1399.77 = 3.22* 10^-5 N

=====

b)

F = qV B = 0

======

c)

F = 3.22* 10^-5 N

======

d)

B = E/ c = 1399.77 / ( 3* 10^8) = 4.666* 10^-6 T

Force exerted on particle

F = q v B = 2.3* 10^-8* 3.7* 10^4* 4.666* 10^-6

F = 3.971* 10^-9 N